• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Roedelbronn

    Log entry time 03:24:08 on June21,2002

    Entry number 86536

    keyword=Target problem

    OK, so "Target problem" is a bit of an understatement... At around 02:25, temperatures on both loops started rising rapidly, and coolant flow to several magnets indicated low. We immediately turned off the beam and ensured the HP heaters were off. The ESR charts indicated that all coolant flows had gone to zero. I called JP, and he advised opening the warm return after contacting MCC. I called MCC and we coordinated going through the warmup procedurein order to place the target in a known condition. Both JT valves were shut per step 5 of the warmup procedure, and the warm return was opened per step 9. MCC shut the cold return per steps 7, 8, 10-13. We did not request the 15K supply valve to be shut, so I assume it is still open (step 13). Warmup procedure complete at 03:12.
    JP doesn't believe the target will be ready for cooldown before the end of shift, but we have no status report from the cryo people or MCC.
    The H2 Vent Open light on the alarm/stutdown box is blinking; no lights on the D2 side